OpenStudy (anonymous):

differentiate y=e^2pix

OpenStudy (cggurumanjunath):

2*pi*e^(2*pi*x) is the final answer very easy question,d/dx(e^(a*x))=a*e^(a*x)

OpenStudy (cggurumanjunath):

ask me any question you want on differentiation i am always there for your help.

OpenStudy (anonymous):

y= e\[y=e ^{2}\pi x\]

OpenStudy (anonymous):

i men y= y prime

OpenStudy (anonymous):

Im correct?

OpenStudy (cggurumanjunath):

no you are wrong,there is slight confusion in the question itself y=e2πx\[y=e ^{2}*\pi*x\] then y'=e ^{2}*\pi\[e ^{2}*\pi\]